OpenStudy (anonymous):

Wind farms are a source of renewable energy found around the world. The power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). If a turbine generates 700kW in a 5 m/s wind, how much power does it generate in a 14 m/s wind? A. 76,832 kw B. 5,488 kW C. 15,366 kW D. 30,733 kW

OpenStudy (anonymous):

\[P = a \times v^3\] Input given values: \[700kW = a \times 5^3\] Solve for a: \[a = 5.6000\] Now use: \[P = a \times v^3\] \[P = 5.6000 \times 14^3\] \[P = 15366.4\] \[15366.4kW\]